Solve for x:
1) 3^x=729 I got 6
2) 6^x-28=-8 I got 10/3
3) 4ln 3x =15 I got 14.17
am I correct?

Answer by fractalier(2101)

Okay, let's see...
1) 3^x = 729 I got 6 ====> CORRECT
2) 6^x - 28 = -8 I got 10/3 ====> not correct. First add 28 to both sides...
6^x = 20
Now take the log of both sides and solve for x...
ln 6^x = ln 20
x = ln 20 / ln 6 and
x = 1.672
3) 4ln 3x =15 I got 14.17 ====> CORRECT. Here divide by four, then exponentiate and solve and we get
ln 3x = 15/4
3x = e^(15/4)
x = e^(15/4) / 3
x = 14.17